OpenStudy (anonymous):

Find the axis of symmetry for this parabola: y = -x2 − 2x − 5 is -3 the axis of symmetry

OpenStudy (anonymous):

Hold on I'll help. Just need a second

OpenStudy (anonymous):

x^2 + 2x - 5 = y Now we need to rearrange x^2 + 2x + ? = y-5

OpenStudy (michele_laino):

the axis of a parabola y = ax^2+bx+c, is given by the subsequent equation:: \[\Large x = - \frac{b}{{2a}}\]

OpenStudy (anonymous):

in this case a = 1, and b = 2 right? @Michele_Laino

OpenStudy (michele_laino):

I think a=-1 and b=-2 since we have to look at the original equation

OpenStudy (anonymous):

oh yes I forgot that there are negative values

OpenStudy (anonymous):

So that means the equation would be\[x= -\frac{ -2 }{ 2(-1)}\]

OpenStudy (anonymous):

Or \[x = -\frac{ -2 }{ -2 }\]

OpenStudy (anonymous):

Or \[x = - \frac{ 1 }{ 1 }\]

OpenStudy (anonymous):

Or x = 1

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

I thin, it is x=-1

OpenStudy (michele_laino):

think*